OpenStudy (anonymous):

a 2150kg truck is travelling along a straight road at a constant speed of 55km/h when the driver removes his foot from the gas. After 21s the trucks speed is 33km/h. What is the magnitude of the average net force acting on the truck during the 21s interval. 626N 1890N 972N 2250N 229N

OpenStudy (snowsurf):

You use the same equations like in the previous problem except you go backwards. You find acceleration then you can find the force since you know the mass of the truck.

OpenStudy (anonymous):

I tried that I did 55(2150)/21 and did not get the right answer

OpenStudy (anonymous):

@snowsurf

OpenStudy (snowsurf):

Lets work this out together. List me the information you are given.

OpenStudy (anonymous):

m=2150kg vi=55km/h vf=33km/h t=21s @snowsurf

OpenStudy (snowsurf):

Great. Now we need an equation that can find us acceleration using the information you have. There are several kinematic equations but one will work. You need an equation that can related vf, vi and t to get a.

OpenStudy (anonymous):

is it the one where v=vo+1/2at ? @snowsurf

OpenStudy (snowsurf):

You have it the correct form just remove the 1/2. Then solve for a.

OpenStudy (snowsurf):

Also want to mention you have to covert their speed from km/h to m/s

OpenStudy (anonymous):

vi-vf/t =55-33/21 =1.05 But I did that in km/h @snowsurf

OpenStudy (snowsurf):

You can use Google calculator to convert it.

OpenStudy (snowsurf):

But you do have the right idea. So use Google to convert and then calculate acceleration a.

OpenStudy (anonymous):

OKay I got 6.11113 @snowsurf

OpenStudy (snowsurf):

Okay. Great that is your a. So a =6.1, we will round it so we don't have so many 1. Now you know acceleration and mass of truck. Now what equation allow you to calculate force?

OpenStudy (anonymous):

F=ma =2150x6.1 =13115N @snowsurf

OpenStudy (snowsurf):

Opps I look at the problem it says average force. My bad. This won't give you the answer. This is only for force at a given instant of time. We are given\[\Delta t\]which is average time. The equation that will relate to this is impulse equation.

OpenStudy (snowsurf):

\[F _{avg}\Delta t = m \Delta v\] That is the impulse equation.

OpenStudy (snowsurf):

Did you calculate it?

OpenStudy (anonymous):

Sorry Open study keeps freezing on me and would the equation be F=mv/t? F=(2150)(6.1)/21

OpenStudy (snowsurf):

That is correct.